There were 4 bike riders.
Paul rode 3.75 km in 15 min.
Lisa rode 4.2 km in 18 min.
Carol rode 3.2 km in 12 min.
Eric rode 5.5 km in 30 min.

Who rode at the greatest rate of speed?
I already had in mind that it was Carol. I found that out when I divided each minutes by the km. This is what I got:
Paul: 0.25 km per minute
Lisa: 0.23 km per minute
Carol: 0.26 km per minute
Eric: 0.18 km per minute
I'm not sure if that was correct what I did, and please help me out and see if you can figure it out for me.
Thank you so much!!!
---

Answer by rothauserc(4718)

This requires the use of the following formula:
distance(d) = rate(r) * time(t) and r = d / t
Paul's rate = 3.75 / 15 = .25 km/minute
Lisa's rate = 4.2 / 18 = .23 km/minute
Carol's rate = 3.2 / 12 = .27 km/minute
Eric's rate = 5.5 / 30 = .18 km/minute
Carol's is the fastest rate